# Quantum Pufferfish Privacy: A Flexible Privacy Framework for Quantum Systems

### v1

@@ Appendix A Proof of Lemma 2
$\sqrt{F(\rho,\widetilde{\rho})}$ $\geq\frac{1}{\operatorname{Tr}[G^{{\dagger}}G\rho]}\left|\langle\psi|_{RA}I_{R}\otimes G_{A}|\psi\rangle_{RA}\right|$ (315)
@@
Appendix B Subadditivity of Smooth Max-Relative Entropy
@@
Given $\delta_{1},\delta_{2}\in\left[0,1\right]$, states $\rho_{1}$ and $\rho_{2}$, and PSD operators $\sigma_{1}$ and $\sigma_{2}$, the following subadditivity relation holds
@@
Appendix C Proof of Theorem 1
For (1): Fix $(\rho^{\mathcal{R}},\rho^{\mathcal{T}})$ in (23), $\mathrm{M}\in\mathcal{M}$. Consider
@@ Appendix C Proof of Theorem 1
where (a) follows due to each $\mathcal{A}_{i}$ being $(\varepsilon,\delta)$-QPP. This relation is true for every $(\rho^{\mathcal{R}},\rho^{\mathcal{T}})$, and so it is true for all such pairs generated from $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$.
@@
because $\mathcal{N}^{\dagger}$ is positive and unital by the assumption $\mathcal{N}$ is a quantum channel. Similarly, we have that $\mathrm{Tr}\!\left[\mathrm{M}^{\prime}\mathcal{N}\left(\mathcal{A}(\rho^{\mathcal{T}})\right)\right]=\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]$, and we conclude that the processed mechanism satisfies $(\varepsilon,\delta)$-QPP with the choice of $\mathcal{M}^{\prime}\subseteq\{\mathrm{M}^{\prime}:\mathcal{N}^{\dagger}(\mathrm{M}^{\prime})\in\mathcal{M}\}$.
@@
and $\mathcal{A}^{\otimes k}(\rho^{\mathcal{T}^{(k)}})$ similarly by replacing $\mathcal{R}$ with $\mathcal{T}$.
@@
Appendix D Proof of Proposition 3
@@
Appendix E Proof of Lemma 4
@@
The proof follows analogously to the classical version of this bound in [70, Proposition 3.3], along with the upper bound for an arbitrary $\mathsf{D}_{\alpha}(\cdot\|\cdot)$ satisfying data processing. Set $\alpha>1$. Then, for such $\mathsf{D}_{\alpha}(\cdot\|\cdot)$, from [65, Equation 4.34], which is obtained by choosing a specific preparation channel, we have that
@@
where the first inequality follows from the inequality in (382), the second from the convexity of the function $x^{\alpha}$ for $\alpha>1$, and the definitions
@@
Together with [70, Lemma B.1], we can further bound (394) from above by $e^{\alpha(\alpha-1)\varepsilon^{2}/2}$. With the substitution in (383) we conclude the proof.

### v2

@@ Appendix B Proof of Lemma 2
$\sqrt{F(\rho,\widetilde{\rho})}$ $\geq\frac{1}{\sqrt{\operatorname{Tr}[G^{{\dagger}}G\rho]}}\left|\langle\psi|_{RA}I_{R}\otimes G_{A}|\psi\rangle_{RA}\right|$ (265)
@@
Appendix C Subadditivity of Smooth Max-Relative Entropy
@@
Given $\delta_{1},\delta_{2}\in\left[0,1\right]$ such that $\delta_{1}+\delta_{2}\leq 1$, states $\rho_{1}$ and $\rho_{2}$, and PSD operators $\sigma_{1}$ and $\sigma_{2}$, the following subadditivity relation holds
@@
Appendix D Proof of Theorem 1
For (1): Fix $(\rho^{\mathcal{R}},\rho^{\mathcal{T}})$ in (23), $\mathrm{M}\in\mathcal{M}$. Consider that
@@ Appendix D Proof of Theorem 1
where (a) follows due to each $\mathcal{A}_{i}$ being $(\varepsilon,\delta)$-QPP. This inequality holds for every $(\rho^{\mathcal{R}},\rho^{\mathcal{T}})$, and so it holds for all such pairs generated from $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$.
@@
because $\mathcal{N}^{\dagger}$ is positive and unital by the assumption that $\mathcal{N}$ is a quantum channel. Similarly, we have that $\mathrm{Tr}\!\left[\mathrm{M}^{\prime}\mathcal{N}\left(\mathcal{A}(\rho^{\mathcal{T}})\right)\right]=\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]$, and we conclude that the processed mechanism satisfies $(\varepsilon,\delta)$-QPP with the choice of $\mathcal{M}^{\prime}\subseteq\{\mathrm{M}^{\prime}:\mathcal{N}^{\dagger}(\mathrm{M}^{\prime})\in\mathcal{M}\}$.
@@
and $\mathcal{A}^{(k)}(\rho^{\mathcal{T}^{(k)}})$ similarly by replacing $\mathcal{R}$ with $\mathcal{T}$.
@@
Appendix E Proof of Proposition 3
@@
Appendix F Composability with Classically Correlated States
In Property 3 of Theorem 1 and Proposition 3, we considered the case in which two mechanisms, composed either in parallel or adaptively, receive independent inputs (i.e., the input being $\rho^{X_{1}}\otimes\rho^{X_{2}}$ where $X_{i}\sim P_{X}\in\Theta$ for $i=\{1,2\}$, which are chosen independently). We now focus on the setting in which the inputs are classically correlated. The input is chosen as a separable state of the form
where $q$ represents a probability distribution with $q(z)\geq 0$ and $\sum_{z\in\mathcal{Z}}q(z)=1$, and $\omega^{z}$ and $\tau^{z}$ are quantum states for all $z\in\mathcal{Z}$ [^11] Note that (332) covers the case of having input states of the form $\sigma_{I}\coloneqq\sum_{(x,y)\in\mathcal{X}\times\mathcal{Y}}q(x,y)\ \omega^{x}\otimes\tau^{y}$ where for all $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ $\omega^{x}$ and $\tau^{y}$ are states, by considering $z$ to be an index for multiple variables, i.e., setting $z=(x,y)$.. One special case of interest is as follows:
$\sigma^{1}_{I}\coloneqq\sum_{z\in\mathcal{Z}}q(z)\,\omega_{1}^{z}\otimes\tau_{1}^{z}\quad\textnormal{and}\quad\sigma^{2}_{I}\coloneqq\sum_{z\in\mathcal{Z}}q(z)\,\omega_{2}^{z}\otimes\tau_{2}^{z},$ (334)
Corollary 4 (Composability of flexible QDP).
  1. 1.
  2. 2.
Proof:
Remark 26 (Extensions beyond flexible QDP).
$\Theta\subseteq\left\{P_{X}\in\mathcal{P}(\mathcal{X}):\begin{array}[]{cc}\forall\ (\mathcal{R},\mathcal{T})\in\mathcal{Q},\ \exists\ x,x^{\prime}\in\mathcal{X}&\\ \textnormal{s.t.}\ q_{\mathcal{R}}(x)=q_{\mathcal{T}}(x^{\prime})=1\end{array}\right\}$ (337)
Appendix G Characterizing Optimal Privacy-Utility Tradeoff
In this Appendix, we focus on identifying the optimal utility that can be obtained by applying an $(\varepsilon,\delta)$-QPP mechanism. Here, we first focus on the setting in which $\mathcal{Q}=\{(\mathcal{R}_{1},\mathcal{R}_{2}),(\mathcal{R}_{2},\mathcal{R}_{1})\}$, ¯=$\bar{\mathcal{M}}=\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\}$, and $\Theta=\{P_{X}\}$, but the following ideas can be extended to the case when $\mathcal{Q}$ is an arbitrary finite set and $\Theta$ includes a finite number of probability distributions. However, the computational complexity involved in identifying the optimal utility increases with the cardinality of the set $\mathcal{Q}$ and $\Theta$, due to the addition of more constraints to the optimization problem.
To incorporate privacy requirements, we use the equivalent formulation of QPP via the DL divergence presented in Proposition 1. To this end, first, we employ the SDP formulated in Lemma 1 to compute the relevant DL divergence and then use that in the optimization of utility. We showcase the use of this SDP in characterizing optimal utility next.
Proposition 14 (Optimal utility for fixed privacy constraints).
$\mathsf{U}\!\left(\varepsilon,\delta,\mathcal{R}_{1},\mathcal{R}_{2}\right)\coloneqq 1-\\ \inf_{\begin{subarray}{c}\mu\geq 0\\ Z_{AD}\geq 0\\ \Gamma_{CD}^{\mathcal{B}}\geq 0\\ \Gamma_{AC}^{\mathcal{A}}\geq 0\\ \lambda_{1}\geq 0,Y_{1}\geq 0\\ \lambda_{2}\geq 0,Y_{2}\geq 0\end{subarray}}\left\{\begin{array}[c]{c}\mu:\\ Z_{AD}\geq\Gamma_{AD}-\mathrm{Tr}_{C}\!\left[\Gamma_{CD}^{\mathcal{B}}\mathrm{T}_{C}(\Gamma_{AC}^{\mathcal{A}})\right],\\ \mu I_{A}\geq\mathrm{Tr}_{D}\!\left[Z_{AD}\right],\\ \mathrm{Tr}_{D}\!\left[\Gamma_{CD}^{\mathcal{B}}\right]=I_{C},\\ \mathrm{Tr}_{D}\!\left[\Gamma_{AC}^{\mathcal{A}}\right]=I_{A},\\ \ln(\lambda_{1})\leq\varepsilon,\\ \mathrm{Tr}\!\left[Y_{1}\right]\leq\delta,\\ Y_{1}\geq\mathrm{Tr}_{A}\!\left[\big{(}\mathrm{T}(\rho^{\mathcal{R}_{1}})\otimes I_{C}\big{)}\Gamma_{AC}^{\mathcal{A}}\right]\\ \quad-\lambda_{1}\mathrm{Tr}_{A}\!\left[(\mathrm{T}(\rho^{\mathcal{R}_{2}})\otimes I_{C})\Gamma_{AC}^{\mathcal{A}}\right],\\ \ln(\lambda_{2})\leq\varepsilon,\\ \mathrm{Tr}\!\left[Y_{2}\right]\leq\delta,\\ Y_{2}\geq\mathrm{Tr}_{A}\!\left[(\mathrm{T}\big{(}\rho^{\mathcal{R}_{2}}\big{)}\otimes I_{C})\Gamma_{AC}^{\mathcal{A}}\right]\\ \quad-\lambda_{2}\mathrm{Tr}_{A}\!\left[\big{(}\mathrm{T}(\rho^{\mathcal{R}_{1}})\otimes I_{C}\big{)}\Gamma_{AC}^{\mathcal{A}}\right]\end{array}\right\}.$ (338)
Proof:
The proof follows from the SDP formulation of the $\gamma$-utility given in Proposition 6, and the privacy constraints (i.e., $\max\{\overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho^{\mathcal{R}_{1}})\|\mathcal{A}(\rho^{\mathcal{R}_{2}})\right),\overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho^{\mathcal{R}_{2}})\|\mathcal{A}(\rho^{\mathcal{R}_{1}})\right)\}\leq\varepsilon$) imposed through the SDP formulation of DL divergence presented in Lemma 1. We also used the fact that for a superoperator $\mathcal{A}$ from system $A$ to $C$, the following equality holds [37, Eq. (3.2.14)]
can also be incorporated to compute the optimum utility.
Remark 28 (Optimal privacy parameters for fixed utility).
$\lambda^{\star}_{1}(\mathcal{A},\gamma,\delta)\coloneqq\\ \inf_{\begin{subarray}{c}\lambda\geq 0\\ Z_{AD}\geq 0\\ \Gamma_{CD}^{\mathcal{B}}\geq 0\\ Y_{1}\geq 0\end{subarray}}\left\{\begin{array}[c]{c}\lambda:\\ Z_{AD}\geq\Gamma_{AD}-\mathrm{Tr}_{C}\!\left[\Gamma_{CD}^{\mathcal{B}}\mathrm{T}_{C}(\Gamma_{AC}^{\mathcal{A}})\right],\\ (1-\gamma)I_{A}\geq\mathrm{Tr}_{D}\!\left[Z_{AD}\right],\\ \mathrm{Tr}_{D}\!\left[\Gamma_{CD}^{\mathcal{B}}\right]=I_{C},\\ \mathrm{Tr}\!\left[Y_{1}\right]\leq\delta,\\ Y_{1}\geq\mathrm{Tr}_{A}\!\left[\big{(}\mathrm{T}(\rho^{\mathcal{R}_{1}})\otimes I_{C}\big{)}\Gamma_{AC}^{\mathcal{A}}\right]\\ \quad-\lambda\mathrm{Tr}_{A}\!\left[(\mathrm{T}(\rho^{\mathcal{R}_{2}})\otimes I_{C})\Gamma_{AC}^{\mathcal{A}}\right]\end{array}\right\}.$ (341)
$\varepsilon^{\star}(\mathcal{A},\gamma,\delta)\coloneqq\ln\!\left(\max\{\lambda^{\star}_{1}(\mathcal{A},\gamma,\delta),\lambda^{\star}_{2}(\mathcal{A},\gamma,\delta)\}\right).$ (342)
Appendix H Proof of Lemma 4
@@
The proof follows analogously to the classical version of this bound in [82, Proposition 3.3], along with the upper bound for an arbitrary $\mathsf{D}_{\alpha}(\cdot\|\cdot)$ satisfying data processing. Set $\alpha>1$. Then, for such $\mathsf{D}_{\alpha}(\cdot\|\cdot)$, from [79, Equation 4.34], which is obtained by choosing a specific preparation channel, we have that
@@
where the first inequality follows from the inequality in (343), the second from the convexity of the function $x\mapsto x^{\alpha}$ for $\alpha>1$, and the definitions
@@
Together with [82, Lemma B.1] and the assumption that $\alpha\varepsilon\leq 2$, we can further bound (355) from above by $e^{\alpha(\alpha-1)\varepsilon^{2}/2}$. With the substitution in (344) we conclude the proof.
Appendix I Proof of Lemma 5
With the choice for $\delta^{\prime}$ as in the Lemma statement (i.e., (206)), we have $\delta-\delta^{\prime}=(1-\delta)(e^{\varepsilon^{\prime}}-e^{\varepsilon})/(e^{\varepsilon}+1)$. Plugging this in, we find that
$\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{R}})\right]$
By choosing the measurement operator $\mathrm{I}-\mathrm{M}$, we also have by assumption that
Rewriting (359), we arrive at
Similar to the previous manipulations, we get
$\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{R}})\right]\leq e^{\varepsilon^{\prime}}\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]+\delta^{\prime}\\ \quad\quad+(e^{\varepsilon^{\prime}}-e^{-\varepsilon})\left(-\mathrm{Tr}\!\left[\mathrm{M}\mathcal{A}(\rho^{\mathcal{T}})\right]+\frac{1-\delta}{e^{\varepsilon}+1}\right).$ (361)
A similar inequality holds for every $(\mathcal{R},\mathcal{T})\in\mathcal{Q}$ and $P_{X}\in\Theta$. Thus, the desired implication has been proved.
Appendix J Bounds on Holevo Information from QDP and QLDP
Proposition 15 (Bounds on Holevo information due to QLDP).
Furthermore, if $\mathcal{A}$ is $(\varepsilon,\delta)$-QLDP, then
is the binary Shannon entropy in nats.
Proof:
The following proof ideas are inspired by the works on mutual information based classical DP and PP presented in [11, Theorem 1] and [10, Theorem 1], respectively. For the proof of the first part, consider that
$\mathsf{I}\!\left(X;B\right)_{\sigma}$
$\leq\min\!\left\{\varepsilon,\frac{\varepsilon^{2}}{2}\right\},$ (369)
For the second part, consider that
$\leq\delta^{\prime}\ln(d-1)+\mathsf{h}(\delta^{\prime}),$ (371)
where
$\bar{\rho}=\sum_{x\in\mathcal{X}}P_{X}(x)\rho^{x},$ (372)
and the final inequality follows from the fact that $\left\|\mathcal{A}(\bar{\rho})-\mathcal{A}(\rho^{x})\right\|_{1}\leq 2\delta^{\prime}$ due to $\mathcal{A}$ satisfying $(\varepsilon,\delta)$-QPP along with Proposition 10. Then, to arrive at the final expression, we use the entropy continuity result known as the Fannes–Audenaert Inequality [89]:
concluding the proof. ∎
Proposition 16 (Bounds on mutual information due to QDP).
Proof:
$\omega^{x}$ $\coloneqq\rho_{i}^{x}\otimes\rho^{n\backslash i}.$ (379)
For the first part, it thus follows that
For the second part, consider that
$\mathsf{I}(X_{i};B|X^{n\backslash i}=z^{n\backslash i})_{\sigma}$
$\leq\delta^{\prime}\ln(d-1)+\mathsf{h}(\delta^{\prime}),$ (382)
where the last inequality holds because $\left\|\mathcal{A}(\omega)-\mathcal{A}(\omega^{x})\right\|_{1}\leq 2\delta^{\prime}$ with $\mathcal{A}$ being $(\varepsilon,\delta)$-QDP for the pair $\sigma\sim\sigma^{x}$, and again applying the continuity result for quantum entropy, as in the proof of Proposition 15. ∎
By the data processing inequality for mutual information, we have $\mathsf{I}(X_{i};Y|X^{n\backslash i})\leq\varepsilon^{\prime}\coloneqq\min\{\varepsilon,\varepsilon^{2}/2\}$. This showcases that the setup of Proposition 16 satisfies $\varepsilon^{\prime}$-mutual information differential privacy, as proposed in [11], where a randomized mechanism $A:\mathcal{X}^{n\times k}\to\mathcal{Y}$ is defined to be $\varepsilon$-mutual information differentially private if